OpenStudy (johnjules2000):

Solve for x: X - + 3 = -4 2

OpenStudy (anonymous):

Would you please rewrite it? It says: X - other stuff 2 is that x + 3 = -4? or -2x + 3 = -4?

OpenStudy (johnjules2000):

it is x divided by 2 plus 3 equals negative 4

OpenStudy (anonymous):

Oh ok: x/2 + 3 = -4 // Subtract 3 from both sides x/2 = -7 // Multiply both sides by 2 x = -14 And check: -14/2 + 3 = -7 + 3 = -4 v/ It works!

OpenStudy (johnjules2000):

I have to show my work. I know the answer is -14, just not sure how to get there.

OpenStudy (johnjules2000):

Thank you Misha!!

OpenStudy (anonymous):

:)